# Higher-order Sobolev and Rellich inequalities via iterated Talenti’s principle

### v1

@@ 1. Introduction
The common key ingredient in establishing the aforementioned results, is the following sharp isoperimetric inequality: If $\Omega\subset\mathbb{R}^{n}$ is a bounded set, then
@@
The goal of the paper is to show that if an isoperimetric inequality holds on a Riemannian manifold, then various Euclidean functional inequalities extend naturally to that setting. Throughout the paper, let $(M,g)$ be an $n$-dimensional Riemannian manifold, and suppose that there exists a constant $c_{g}>0$ such that for every bounded $\Omega\subset M$ one has
@@
Note that this assumption is reasonable according to Brendle’s isoperimetric inequality [5] on non-negatively curved spaces and to Croke’s isoperimetric inequality [8] on non-positively curved spaces, see §2.1.
@@
$|D_{g}^{m}u|\stackrel{{\scriptstyle\textup{def}}}{{=}}\begin{cases}|\Delta_{g}^{k}u|,&\text{if }m=2k,\\ |\nabla_{g}(\Delta_{g}^{k}u)|,&\text{if }m=2k+1.\end{cases}$
@@
Note that $S_{m,p}$ is the greatest constant for which inequality (2) holds when $M=\mathbb{R}^{n}$ (and $c_{g}=1$). If $p=2$, according to Cotsiolis and Tavoularis [7], its exact value is given by
@@
Note that $R_{m,p}$ is the greatest constant for which inequality (3) hold when $M=\mathbb{R}^{n}$ (and $c_{g}=1$). According to Mitidieri [19], its exact value is given by
@@
Our third result is an extension of the previous theorem, which allows for discussing ‘improved’ Rellich inequalities.
@@
The paper is structured as follows. In §2 we present an iterated Talenti’s principle. First, in §2.1 we recall various isoperimetric inequalities. Next, in §2.2 we display the definition and relevant properties of Schwarz symmetrization in the Riemannian setting. Finally, in §2.3 we formulate the principle together with its iterated version (see Theorem 2.3). In §3 we prove Theorem 1.1, that is we establish higher-order Sobolev inequalities on Riemannian manifolds supporting the inequality (1). Moreover, in Proposition 3.1 in the particular case of non-positive Ricci curvature, Euclidean volume growth, and $m=p=2$, we present a sharpness result. In §4 we prove Theorem 1.2, that is we establish higher-order Rellich inequalities on Riemannian manifolds supporting the inequality (1). Here we also prove Theorem 1.3, which allows for establishing improved higher-order Rellich-type inequalities. We conclude the paper with various applications of Theorem 1.3.
@@ 2.1. Isoperimetric inequalities
where $\operatorname{vol}_{g}$ stands for the Riemannian volume induced by $g$; $B_{g}(x,r)$ denotes a geodesic ball in $M$ centered at $x\in M$ having radius $r$; while $\omega_{n}$ denotes the volume of the $n$-dimensional Euclidean unit ball. On the one hand, we note that the above limit is independent of the choice of $x\in M$. On the other hand, according to the Bishop–Gromov comparison theorem (see e.g., Gallot, Hulin, and Lafontaine [9, Theorem 3.101]), one has $\mathsf{AVR}_{g}\in[0,1]$. The typical assumption in this setting is that the manifold $(M,g)$ has Euclidean volume growth, that is
@@
Under this assumption the following sharp isoperimetric inequality holds (see Brendle [5, Corollary 1.3]): For every bounded domain $\Omega\subset M$ one has
@@
Case 3: $\mathbf{Ric}_{g}\geq k\in\mathbb{R}$ and positive lower bound on volume of unit balls. According to Hebey [13, Lemma 3.2], suppose that $(M,g)$ is a complete $n$-dimensional Riemannian manifold with $\mathbf{Ric}_{g}\geq k$ for some $k\in\mathbb{R}$, and there exists $v>0$ such that
$\operatorname{vol}_{g}(B_{g}(x,1))\geq v,\quad\forall x\in M.$
Then there exist $c_{g},\eta>0$ depending only on $n$, $k$, and $v$ such that for every open $\Omega\subset M$ with smooth boundary, compact closure, and $\operatorname{vol}_{g}(\Omega)\leq\eta$, the isoperimetric inequality (1) holds.
Case 4: Bounded elliptic Kato constant As in Impera, Rimoldi, and Veronelli [14, Theorem 1.1], let $(M,g)$ be a complete non-parabolic Riemannian manifold with
@@ 2.2. Schwarz symmetrization
If $\Omega\subset M$ is a bounded domain, its symmetrization, denoted by $\Omega^{\star}$, is an Euclidean ball, centered at the origin satisfying
@@
In the sequel, we present three fundamental properties of symmetrized functions, which are exploited in the rest of the paper. First, the Cavalieri principle (see e.g. Lieb and Loss [17, Theorem 3.3/(iv)]) states that if $u\in L^{q}(\Omega)$ for some $q\in(0,\infty)$, then
@@
Next, according to the Pólya–Szegő inequality (see e.g. Balogh and Kristály [2, Proposition 3.1]), if $|\nabla_{g}u|\in L^{p}(\Omega)$ for some $p>1$ and some smooth function $u$, then one has
@@
Finally, due to the Hardy–Littlewood inequality (see e.g. Lieb and Loss [17, Theorem 3.4]), if $x_{0}\in\Omega$ and $\rho(x)=d_{g}(x_{0},x)$, $u\in L^{p}(\Omega)$ for some $p>1$, and $h\colon[0,\infty)\to[0,\infty)$ is a non-increasing function, then
@@
We note that some authors condiser an alternative symmetrization (tipically denoted by ). In this case, $\Omega^{\sharp}$ is an Euclidean ball satisfying $\operatorname{vol}_{g}(\Omega)=(c_{g})^{n}\cdot\operatorname{vol}_{g_{0}}(\Omega^{\sharp})$ instead of (6), and inequality (7) becomes $\operatorname{per}_{g}(\partial\Omega)\geq(c_{g})^{n}\cdot\operatorname{per}_{g_{0}}(\partial\Omega^{\sharp})$. At first sight, the same exponent of $c_{g}$ in the latter relations may seem to have computational advantage. However, when proving the alternative versions of the fundamental properties (9), (10) & (11), one needs to compare sets with the same volume, and then to perform changes of variables of the form $y=c_{g}\cdot x$. Using the $\Omega^{\star}$ symmetrization, the proofs can be done is the same spirit as their Euclidean counterparts, and no change of variables is required.
@@ 2.3. An iterated Talenti’s principle
The following result is a Talenti’s principle on Riemannian manifolds supporting the isoperimetric inequality (1). The statement is formulated using the notations of the previous section. The proof is analogous to the main result of Chen and Li [6, Theorem 1.1], which expoits the isoperimetric inequality (4).
@@
Let $(M,g)$ is complete, non-compact Riemannian manifold supporting inequality (1). Assume that $\Omega\subset M$ is a bounded domain with smooth boundary and let $\Omega^{\star}$ denote its symmetrization. Suppose that $u$ and $v$ are weak solutions of the problems
@@
Let $(M,g)$ be a complete, non-compact Riemannian manifold supporting inequality (1). Assume that $\Omega\subset M$ is a bounded domain with smooth boundary and let $\Omega^{\star}$ denote its symmetrization. Let $k\geq 1$ be an integer, and suppose that $u$ and $v$ are weak solutions of the problems
@@
$\begin{cases}(-\Delta_{g})^{k-1}\widetilde{u}=f,&\text{in }\Omega,\\ \widetilde{u}=\Delta_{g}\widetilde{u}=\dots=\Delta_{g}^{k-2}\widetilde{u}=0,&\text{on }\partial\Omega,\end{cases}\quad\text{and}\quad\begin{cases}(-\Delta_{g_{0}}\widetilde{v})^{k-1}=f^{\star},&\text{in }\Omega^{\star},\\ \widetilde{v}=\Delta_{g_{0}}\widetilde{v}=\dots=\Delta_{g_{0}}^{k-1}\widetilde{v}=0,&\text{on }\partial\Omega^{\star}.\end{cases}$
@@
On the other hand, we also have $\widetilde{u}=0$ and $\widetilde{v}=0$ on $\partial\Omega$ and $\partial\Omega^{\star}$, respectively, thus we obtain
@@
Thus, combining the Euclidean maximum principle for $\Delta$ with another application of Theorem 2.2, implies the inequality (16).
@@
We note that the non-iterated version of Talenti’s principle is available on Finsler manifolds with non-negative weighted Ricci curvature (see Mester [18]) and on the so-called $\mathsf{RCD}(0,N)$ spaces (see Wu [21]). Depending on the way of symmetrization (cf. Remark 2.1), inequality (13) may also be written in the form $u^{\sharp}\leq v$, which is the case in [6] & [18]. For simplicity, our results are presented on Riemannian manifolds; however, by following their line of thought, one may extend them to the aforementioned settings, taking into account the technicalities of corresponding spaces.
@@ 3. Sobolev inequalities
The goal of this section is twofold. First, we prove Theorem 1.1, that is, we establish higher-order Sobolev inequalities on Riemannian manifolds supporting inequality (1). Next, we present a sharpness result concerning a particular Sobolev inequality, see Proposition 3.1.
@@
Since $u\in C_{0}^{\infty}(M)$, there exists a bounded domain $\Omega\in M$, such that $u$ (and its derivatives) vanish outside of $\Omega$. Thus, it is enough to prove the Sobolev inequality (2) on an arbitrary bounded domain $\Omega\subset M$. Without loss of generality, one may assume that $u>0$. We distinguish two cases.
@@
$(c_{g})^{2kp}\cdot S_{m,p}\cdot\left(\int_{\Omega^{\star}}(u^{\star})^{p^{*}}\,\mathrm{d}v_{g_{0}}\right)^{\frac{p}{p^{*}}}=(c_{g})^{mp}\cdot S_{m,p}\cdot\left(\int_{\Omega}u^{p^{*}}\,\mathrm{d}v_{g}\right)^{\frac{p}{p^{*}}},$
@@
$\geq(c_{g})^{p}\cdot\int_{\Omega^{\star}}(\nabla_{g_{0}}f^{\star})^{p}\,\mathrm{d}v_{g_{0}}=(c_{g})^{p}\cdot\int_{\Omega^{\star}}|\nabla_{g_{0}}(\Delta_{g_{0}}^{k}v)|^{p}\,\mathrm{d}v_{g_{0}}.$
@@
$=(c_{g})^{mp}\cdot S_{m,p}\cdot\left(\int_{\Omega}u^{p^{*}}\,\mathrm{d}v_{g}\right)^{\frac{p}{p^{*}}},$
@@
The proof can be done in the spirit of Kristály [16, Proposition 2.3 & Theorem 1.1] as follows.
@@
On the other hand, by assumption and the Laplacian comparison theorem applied for the case $\mathbf{Ric}_{g}\geq 0$ (see e.g. Gallot, Hulin, and Lafontaine [9, Theorem 3.101]) we also have
@@
$\mathcal{H}(\lambda,s)\stackrel{{\scriptstyle\textup{def}}}{{=}}\int_{M}(\lambda+d_{g}^{2})^{-n}\,\mathrm{d}v_{g},\quad\lambda,s>0,$
@@
(see e.g. Kristály [16, Proposition 2.3/(ii)]) and the following property of the gamma function:

### v2

@@ 1. Introduction
The common key ingredient in establishing the aforementioned results, is the following sharp isoperimetric inequality: If $\Omega\subset\mathbb{R}^{n}$ is a regular domain, that is a bounded open set with smooth boundary, then
@@
The goal of the paper is to show that if an isoperimetric inequality holds on a Riemannian manifold, then various Euclidean functional inequalities extend naturally to that setting. Throughout the paper, let $(M,g)$ be an $n$-dimensional Riemannian manifold, and suppose that there exists a constant $c_{g}>0$ such that for every regular domain $\Omega\subset M$ one has
@@
Geometric conditions that guarantee the validity of the inequality (1) with an explicit value of $c_{g}$ are the following (we refer to §2.1 for more details):
  1. (1)
    non-negative Ricci curvature and Euclidean volume growth, see Brendle [6];
  2. (2)
    non-positive sectional curvature, see Croke [12];
  3. (3)
    non-parabolicity and bounded elliptic Kato constant, see Impera, Rimoldi, and Veronelli [21].
@@
$|D_{g}^{m}u|\stackrel{{\scriptstyle\textup{def}}}{{=}}\begin{cases}|\Delta_{g}^{k}u|,&\text{if }m=2k,\\ |\nabla_{g}(\Delta_{g}^{k}u)|,&\text{if }m=2k+1,\end{cases}$
where $\Delta^{k}_{g}u$ means that the Laplace–Beltrami operator of $(M,g)$ is applied $k$-times to $u$, while $\nabla_{g}$ stands for the Riemannian gradient.
@@
Note that if $p=1$ (and $m=1$), then the inequality (2) becomes the so-called $L^{1}$-Sobolev inequality, which turns out to be equivalent to the isoperimetric inequality (1). According to Talenti [29], “if $p=1$ the Sobolev inequality behaves in a slightly different manner”, and the determination of the best constant becomes problematic, see e.g. Cassani, Ruf, and Tarsi [8]. In this paper, we focus on the case when $p>1$.
The investigation of the best constant in the inequality (2) in the Euclidean setting (when $m=1$) originates from the pioneering works of Aubin [2] and Talenti [29]. For general $m$ and $p=2$ it is given by
@@
see Cotsiolis and Tavoularis [10].
The sharp version of the inequality (2) on Cartan–Hadamard manifolds (i.e., complete, simply connected Riemannian manifolds having non-positive sectional curvature) is also provided by Aubin, assuming the validity of the so-called Cartan–Hadamard conjecture (that is the sharp isoperimetric inequality on the aforementioned setting); see also Hebey [20, Chapter 8]. Concerning spaces with non-positive Ricci curvature, Ledoux [25] showed that attaining the Euclidean optimal constant in a Sobolev inequality forces $(M,g)$ to be isometric to the Euclidean space. The sharp version of the inequality (2) is established by Brendle [6] using ABP method, and Kristály [24] using optimal mass transport theory.
Concerning the case when $m=2$ and $p=2$, Barbosa and Kristály [4] proved a rigidity result for the corresponding Sobolev inequality in the same spirit as Ledoux. In Theorem 3.1 we explicitly determine (under a mild geometric assumption, see Remark 3.2) the sharp constant for the aforementioned inequality. The issue of sharpness in the general case ($m>2$ and $p>1$) is challenging even in the Euclidean setting.
@@
The investigation of the best constant in the inequality (3) in the Euclidean setting is due to Mitidieri [28], the exact value is given by
@@
Since Hardy–Rellich inequalities typically does not have extremizers, they admit various improvements. For a comprehensive discussion on them we refer to the monograph by Ghoussoub and Moradifam [19], we also refer to Kajántó, Kristály, Peter, and Zhao [22] for an alternative approach to Hardy-type inequalities. The sharpness of inequality (3) can be established also on Cartan–Hadamard manifolds, see Farkas, Kajántó, and Kristály [13]. On spaces with non-positive Ricci curvature the issue of sharpness is still open.
Our third result is an extension of the previous theorem, which allows for discussing improved Rellich inequalities.
@@
The paper is structured as follows. In §2 we present an iterated Talenti’s principle. First, in §2.1 we recall various isoperimetric inequalities. Next, in §2.2 we display the definition and relevant properties of Schwarz symmetrization in the Riemannian setting. Finally, in §2.3 we formulate the principle together with its iterated version (see Theorem 2.4). In §3 we prove Theorem 1.1, that is we establish higher-order Sobolev inequalities on Riemannian manifolds supporting the inequality (1). Moreover, in Theorem 3.1 in the particular case of non-positive Ricci curvature, Euclidean volume growth, and $m=p=2$, we present a sharpness result. In §4 we prove Theorem 1.2, that is we establish higher-order Rellich inequalities on Riemannian manifolds supporting the inequality (1). Here we also prove Theorem 1.3, which allows for establishing improved higher-order Rellich-type inequalities. We conclude the paper with various applications of Theorem 1.3.
@@ 2.1. Isoperimetric inequalities
where $\operatorname{vol}_{g}$ stands for the Riemannian volume induced by $g$; $B_{g}(x,r)$ denotes a geodesic ball in $M$ centered at $x\in M$ having radius $r$; while $\omega_{n}$ denotes the volume of the $n$-dimensional Euclidean unit ball. On the one hand, we note that the above limit is independent of the choice of $x\in M$. On the other hand, according to the Bishop–Gromov comparison theorem (see e.g., Gallot, Hulin, and Lafontaine [15, Theorem 3.101]), one has $\mathsf{AVR}_{g}\in[0,1]$. The typical assumption in this setting is that the manifold $(M,g)$ has Euclidean volume growth, that is
@@
Under this assumption the following sharp isoperimetric inequality holds (see Brendle [6, Corollary 1.3]): For every regular domain $\Omega\subset M$ one has
@@
Case 3: Bounded elliptic Kato constant. As in Impera, Rimoldi, and Veronelli [21, Theorem 1.1], let $(M,g)$ be a complete non-parabolic Riemannian manifold with
@@
Remark 2.1.
We note that there exist various results ensuring the validity of the isoperimetric inequality (1), without specifying the explicit value of the constant $c_{g}$. For example, according to Carron [7], Coulhon and Ledoux [11], and Varopoulos [31], if $(M,g)$ is a complete Riemannian manifold with infinite volume, $n\geq 3$, $\mathbf{Ric}_{g}\geq 0$, then inequality (1) holds for some $c_{g}>0$ if and only if there exists $K>0$ such that for every $x\in M$ and $t>0$ one has
$\operatorname{vol}_{g}(\{y\in M:G_{x}(y)>t\})\leq K\cdot t^{-n/(n-2)},$
where $G_{x}$ is the positive minimal Green function of pole $x$.
@@ 2.2. Schwarz symmetrization
If $\Omega\subset M$ is a regular domain, its symmetrization, denoted by $\Omega^{\star}$, is an Euclidean ball, centered at the origin satisfying
@@
In the sequel, we present three fundamental properties of symmetrized functions, which are exploited in the rest of the paper. First, the Cavalieri principle (see e.g. Lieb and Loss [26, Theorem 3.3/(iv)]) states that if $u\in L^{q}(\Omega)$ for some $q\in(0,\infty)$, then
@@
Next, according to the Pólya–Szegő inequality (see e.g. Balogh and Kristály [3, Proposition 3.1]), if $|\nabla_{g}u|\in L^{p}(\Omega)$ for some $p>1$ and some smooth function $u$, then one has
@@
Finally, due to the Hardy–Littlewood inequality (see e.g. Lieb and Loss [26, Theorem 3.4]), if $x_{0}\in\Omega$ and $\rho(x)=d_{g}(x_{0},x)$, $u\in L^{p}(\Omega)$ for some $p>1$, and $h\colon[0,\infty)\to[0,\infty)$ is a non-increasing function, then
@@
We note that some authors condiser an alternative symmetrization: $\Omega^{\sharp}$ is an Euclidean ball satisfying $\operatorname{vol}_{g}(\Omega)=(c_{g})^{n}\cdot\operatorname{vol}_{g_{0}}(\Omega^{\sharp})$ instead of (6), and inequality (7) becomes $\operatorname{per}_{g}(\partial\Omega)\geq(c_{g})^{n}\cdot\operatorname{per}_{g_{0}}(\partial\Omega^{\sharp})$. It turns out that when proving the alternative versions of the fundamental properties (9), (10) & (11), one needs to compare sets with the same volume, and then to perform changes of variables of the form $y=c_{g}\cdot x$. Using the $\Omega^{\star}$ symmetrization, the proofs can be done is the same spirit as their Euclidean counterparts, and no change of variables is required.
@@ 2.3. An iterated Talenti’s principle
The following result is a Talenti’s principle on Riemannian manifolds supporting the isoperimetric inequality (1). The statement is formulated using the notations of the previous section. The proof is the same as the one of Chen and Li [9, Theorem 1.1], which exploits the isoperimetric inequality (4), and is therefore omitted.
@@
Let $(M,g)$ is complete, non-compact Riemannian manifold supporting inequality (1). Assume that $\Omega\subset M$ is a regular domain and let $\Omega^{\star}$ denote its symmetrization. Suppose that $u$ and $v$ are weak solutions of the problems
@@
Let $(M,g)$ be a complete, non-compact Riemannian manifold supporting inequality (1). Assume that $\Omega\subset M$ is a regular domain and let $\Omega^{\star}$ denote its symmetrization. Let $k\geq 1$ be an integer, and suppose that $u$ and $v$ are weak solutions of the problems
@@
$\begin{cases}(-\Delta_{g})^{k-1}\widetilde{u}=f,&\text{in }\Omega,\\ \widetilde{u}=\Delta_{g}\widetilde{u}=\dots=\Delta_{g}^{k-2}\widetilde{u}=0,&\text{on }\partial\Omega,\end{cases}\quad\text{and}\quad\begin{cases}(-\Delta_{g_{0}})^{k-1}\widetilde{v}=f^{\star},&\text{in }\Omega^{\star},\\ \widetilde{v}=\Delta_{g_{0}}\widetilde{v}=\dots=\Delta_{g_{0}}^{k-2}\widetilde{v}=0,&\text{on }\partial\Omega^{\star}.\end{cases}$
@@
On the other hand, we also have u=0{u}=0 and v=0{v}=0 on $\partial\Omega$ and $\partial\Omega^{\star}$, respectively, thus we obtain
@@
Finally, let $z$ be a weak solution of
$\begin{cases}-\Delta_{g_{0}}z=\widetilde{u}^{\star},&\text{in }\Omega^{\star},\\ z=0,&\text{on }\partial\Omega^{\star}.\end{cases}$
By Theorem 2.3, we obtain that for a.e. $x\in\Omega$ one has
$(c_{g})^{2}\cdot u^{\star}(x)\leq z(x),$
and the following inequality holds in the weak sense:
$\Delta_{g_{0}}\left((c_{g})^{2(k-1)}\cdot z\right)=(c_{g})^{2(k-1)}\cdot\widetilde{u}^{\star}\leq\widetilde{v}=-\Delta_{g_{0}}v.$
Thus, the Euclidean maximum principle implies the inequality (16).
@@
We note that Fontana, Morpurgo, and Qin [14] used a version of the above iterated Talenti’s principle for Green functions (and Navier boundary conditions) on Riemannian manifolds having Euclidean volume growth. We also note that the non-iterated version of Talenti’s principle is available on Finsler manifolds with non-negative weighted Ricci curvature (see Mester [27]) and on the so-called $\mathsf{RCD}(0,N)$ spaces (see Wu [32]). For simplicity, our results are presented on Riemannian manifolds, but could be extended to the aforementioned settings.
@@ 3. Sobolev inequalities
The goal of this section is twofold. First, we prove Theorem 1.1, that is, we establish higher-order Sobolev inequalities on Riemannian manifolds supporting inequality (1). Next, we present a sharpness result concerning a particular Sobolev inequality, see Theorem 3.1.
@@
Since $u\in C_{0}^{\infty}(M)$, there exists a regular domain $\Omega\in M$, such that $u$ (and its derivatives) vanish outside of $\Omega$. Thus, it is enough to prove the Sobolev inequality (2) on an arbitrary regular domain $\Omega\subset M$. We distinguish two cases.
@@
$(c_{g})^{2kp}\cdot S_{m,p}\cdot\left(\int_{\Omega^{\star}}(u^{\star})^{p^{*}}\,\mathrm{d}v_{g_{0}}\right)^{\frac{p}{p^{*}}}=(c_{g})^{mp}\cdot S_{m,p}\cdot\left(\int_{\Omega}|u|^{p^{*}}\,\mathrm{d}v_{g}\right)^{\frac{p}{p^{*}}},$
@@
$\geq(c_{g})^{p}\cdot\int_{\Omega^{\star}}|\nabla_{g_{0}}f^{\star}|^{p}\,\mathrm{d}v_{g_{0}}=(c_{g})^{p}\cdot\int_{\Omega^{\star}}|\nabla_{g_{0}}(\Delta_{g_{0}}^{k}v)|^{p}\,\mathrm{d}v_{g_{0}}.$
@@
$=(c_{g})^{mp}\cdot S_{m,p}\cdot\left(\int_{\Omega}|u|^{p^{*}}\,\mathrm{d}v_{g}\right)^{\frac{p}{p^{*}}},$
@@
The first part of the proof is given in Barbosa and Kristály [4, Theorem 1.1/(ii)], we reproduce it for the reader’s convenience, the second part can be done in the spirit of Kristály [24, Proposition 2.3 & Theorem 1.1].
@@
On the other hand, by assumption and the Laplacian comparison theorem applied for the case $\mathbf{Ric}_{g}\geq 0$ (see e.g. Gallot, Hulin, and Lafontaine [15, Theorem 3.101]) we also have
@@
$\mathcal{H}(\lambda,s)\stackrel{{\scriptstyle\textup{def}}}{{=}}\int_{M}(\lambda+d_{g}^{2})^{-s}\,\mathrm{d}v_{g},\quad\lambda,s>0,$
@@
(see e.g. Kristály [24, Proposition 2.3/(ii)]) and the following property of the gamma function: